On October 1, 2017, Vaughn, Inc., leased a machine from Fell Leasing Company for five years. The lease requires five annual payments of $10,000 beginning September 30, 2018. Vaughn’s incremental borrowing rate is 11%, and it uses a calendar year for reporting purposes. The machine has a 12-year economic life with zero salvage value. Vaughn correctly classifies the lease as an operating lease under ASU 2016-02. Using (PV of 1, PVAD of 1, and PVOA of 1) (Use the appropriate factor(s) from the tables provided.)

Required:

  1. At what amount should Vaughn record the leased equipment on October 1, 2017? (Round your answer to the nearest whole dollar.)
  2. What is the amount of rent expense that Vaughn should record for the year ended December 31, 2017, and for the year ended December 31, 2018?
  3. How much of the lease liability should be classified as current on December 31, 2017, and December 31, 2018? (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

Part 1)

The amount Vaughn should record the leased equipment on October 1, 2017 is determined as below:

Value of Leased Equipment = Annual Lease Payment*PVOA(Rate, Years)

Here, Annual Lease Payment = 10,000, Rate = 11% and Years = 5

Using PVOA tables, we get,

Value of Leased Equipment = 10,000*PVOA(11%,5) = 10,000*3.6959 = $36,959 (answer for Part 1)

______

Part 2)

The amount of rent expense to be recorded at December 31, 2017 and December 31, 2018 is calculated as follows:

Rent Expense (December 31, 2017) = Annual Lease Payment*3/12 (from October 1, 2017 to December 31, 2017) = 10,000*3/12 = $2,500

Rent Expense (December 31, 2018) = Annual Lease Payment*9/12 (from January 1, 2018 to 30 September 2018) + Annual Lease Payment*3/12 (from October 1, 2018 to December 31, 2018) = 10,000*9/12 + 10,000*3/12 = $10,000

______

Part 3)

The value of lease liability that should be classified as current on December 31, 2017, and December 31, 2018 is arrived with the use of table given below:

Date Lease Payment Interest Expense Current Lease Liability Balance
01-10-2017 36,959.00
31-12-2017 2,500.00 1,016.37 (36,959*3/12*11%) 1,483.63 35,475.37
30-09-2018 7,500.00 2,926.72 (35,475.37*9/12*11%) 4,573.28 30,902.09
31-12-2018 2,500.00 849.81 (30,902.09*3/12*11%) 1,650.19 29,251.90

Based on the calculations in the above table, we can derive the following information:

Current Portion of Lease Liability (31 December, 2017) = $1,483.63

Current Portion of Lease Liability (31 December, 2018) = $1,650.19
Answer #2

To calculate the amount Vaughn should record for the leased equipment on October 1, 2017, we need to find the present value of the lease payments using the incremental borrowing rate of 11%.

Step 1: Calculate the present value of the lease payments: PV of Annuity Due (PVOA) of 1 for 5 periods at 11% = 3.696848 (from the tables)

Present value of lease payments = $10,000 * 3.696848 = $36,968.48

Step 2: Record the leased equipment on October 1, 2017: The leased equipment should be recorded at the present value of the lease payments, which is $36,968.48.

For the amount of rent expense for the year ended December 31, 2017, we need to calculate the interest and reduction of the lease liability.

Step 3: Calculate the interest expense for 2017: Interest expense = Lease liability at the beginning of the year * Incremental borrowing rate Lease liability at the beginning of 2017 = Present value of lease payments = $36,968.48

Interest expense for 2017 = $36,968.48 * 11% = $4,066.53

Step 4: Calculate the reduction of the lease liability for 2017: Reduction of lease liability = Lease payments - Interest expense Lease payments for 2017 = $10,000

Reduction of lease liability for 2017 = $10,000 - $4,066.53 = $5,933.47

Step 5: Calculate the rent expense for 2017: Rent expense for 2017 = Interest expense + Reduction of lease liability Rent expense for 2017 = $4,066.53 + $5,933.47 = $10,000

For the amount of rent expense for the year ended December 31, 2018, we repeat the same calculations using the lease liability at the beginning of 2018.

Lease liability at the beginning of 2018 = Present value of lease payments - Reduction of lease liability for 2017 Lease liability at the beginning of 2018 = $36,968.48 - $5,933.47 = $31,035.01

Step 6: Calculate the interest expense for 2018: Interest expense for 2018 = Lease liability at the beginning of the year * Incremental borrowing rate Interest expense for 2018 = $31,035.01 * 11% = $3,413.86

Step 7: Calculate the reduction of the lease liability for 2018: Reduction of lease liability for 2018 = Lease payments - Interest expense Reduction of lease liability for 2018 = $10,000 - $3,413.86 = $6,586.14

Step 8: Calculate the rent expense for 2018: Rent expense for 2018 = Interest expense + Reduction of lease liability Rent expense for 2018 = $3,413.86 + $6,586.14 = $10,000

For the classification of the lease liability as current on December 31, 2017, and December 31, 2018:

Step 9: Calculate the portion of the lease payments due within one year: Portion of lease payments due within one year = Lease payments Portion of lease payments due within one year = $10,000

As the lease liability is paid annually, the entire lease payments for each year are considered current liabilities.

So, the amount Vaughn should record for the leased equipment on October 1, 2017, is $36,968.48. The rent expense for the year ended December 31, 2017, and December 31, 2018, is $10,000 for each year. The lease liability classified as current on December 31, 2017, and December 31, 2018, is $10,000 for each year.


answered by: mervetokaz